Bond Valuation Exercise

(a)        Compute the price / yield of the following bonds.

(b)        Indicate which bond experiences the biggest price change (in percentage terms) when yields increase by 1%.

Issuer

Bosch

Bank of America

Nissan

Settlement

Today

Today

Today

Coupon

3.50%

2.75%

0%

Frequency coupon payment

1

2

Zero-coupon

Maturity

2 years from now

25 years from now

6 years from now

Face value

€100

$100

€100

Yield

2.2%

3.7%

    6.3%

Price

….

….

….

New Yield

….

….

….

New Price

….

….

….

Percentage change

….

….

….
Answer #1

Answer a and b:

Price of bonds, new yields, new prices and percentage changes are calculated below:

As we observe Bank of America bond experiences biggest percentage change.

Issuer Bosch Bank of America Nissan Settlement Toda Toda Toda Coupon 3.50% 2.75% 0.00% Frequency coupon payment 1 2 Zero-coupon Maturit 2 vears from now 25 vears from now 6 vears from now Face value 100 $100 100 Yield 2.20% 3.70% 6.30% Price 102.52 $84.59 69.31 New Yield 3.20% 4.70% 7.30% New Price 100.57 $71.50 € 65.52 Percentage change 1.90% 15.48% 5.46%

Workings:

Answer a:

Price of Bosch bond:

Face value = €100

Annual coupon amount = 100* 3.50% = $3.50

Years to maturity = 2 years

Yield = 2.2%

Price = PV

PV (rate, nper, pmt, fv, type)

PV (2.2%, 2, -3.50, -100, 0)

= €102.52

Price of Bank of America bond:

Semiannual coupon = $100 * 2.75%/2 =$1.375

Semiannual yield = 3.7%/2 = 1.85%

PV (1.85%, 50, -1.375, -100, 0)

= $84.59

Price of Nissan bond:

PV (6.3%, 6, 0, -100, 0)

= € 69.31

Answer b:

New Yields:

Yield of Bosch =2.2% + 1% = 3.2%

Yield of Bank of America bond = 3.7% + 1% = 4.7%

Yield of Nissan bond =5.3% + 1% = 6.3%

New prices:

Price of Bosch bond = PV (3.2%, 2, -3.50, -100, 0) = €100.57

Price of Bank of America bond = PV (2.35%, 50, -1.375, -100, 0) = $71.50

Price of Nissan bond = PV (7.3%, 6, 0, -100, 0) = € 65.52

Percentage change:

Bosch bond = (100.57 - 102.52) /102.52 = -1.90%

Bank of America bond = (71.50 - 84.59) /84.59 = -15.48%

Nissan bond = (65.52 - 69.31)/69.31 = -5.46%